According to the kinetic molecular theory, if two different gas molecules were at the same temperature and pressure, but one molecule was large and the other was small, which molecule would have the greatest kinetic energy?


What volume of gas would you expect to collect during an experiment if 1.5-moles of carbon dioxide were produced at 35⁰C and 113 kPa?


1
Expert's answer
2021-02-18T07:27:49-0500

According to the kinetic molecular theory, two gases of different molar mass at the same temperature have the same average kinetic energy, though the gas with the smaller molar mass will have a higher root-mean-square speed.


According to the ideal gas law:

pV = nRT

where p - pressure (p = 113 kPa = 1.1 atm), V - volume, n - number of moles (n = 1.5 moles), R - gas constant (0.082 L atm mol-1 K-1), T - temperature (T=  35⁰C = 308.15 K).

from here:

V = nRT / p = 1.5 mol × 0.082 L atm mol-1 K-1 × 308.15 K / 1.1 atm = 34.5 L


Answer: 34.5 L